Consider the following premerger information about a bidding firm (Firm B) and a target firm (Firm T). Assume that both firms have no debt outstanding. Firm B Firm T Shares outstanding 6,000 1,200 Price per share $ 47 $ 17 Firm B has estimated that the value of the synergistic benefits from acquiring Firm T is $9,500. Firm T can be acquired for $19 per share in cash or by exchange of stock wherein B offers one of its share for every two of T's share.

a. Are the shareholders of Firm T better off with the cash offer or the stock offer?

b. At what exchange ratio of B shares to T shares would the shareholders in T be indifferent between the two offers? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

a. Given information,

Firm B Firm T
No. of Outstanding shares 6000 1200
Price per share 47 17

Cash offer: Total payment = 19*1200 = 22,800

Stock offer: Total value of Firm B shares that will be transferred to shareholders of B = (1200/2)*47 = 28,200.

We can clearly see that the stock offer is a better proposition to shareholders of Firm T as they are getting higher value.

If we also take into account the synergistic benefits that will come to Firm B with acquisition of the target Firm T, the price per share of Firm B can be expected to go further up.

Hence, the shareholders of firm T is better off with the stock offer.

b. To answer this question, we need to equate the cash offer and stock offer amount.

19*1200 = 47*1200*(1/Exchange ratio)

Exchange ratio = 47/19 = 2.4737